show that the function f(n) = |n^2 sin n| is in neither O(n) nor \Omega (n)
Answer #1

Answer:

f(n) = | n^2sinn |

Let we assume that f(n) = n^2 for even and f(n) = 0 for odd

Now , say f(n) = O(n) , it will be less than c *n * c*n , where c is constant and c > 0 and similarly f(n) = Omega(n) is not possible also because it will be less than c*n . So both the cases are not possible.